OpenStudy (anonymous):

Any number in the form of a ± bi , where a and b are real numbers and b 0 is considered a pure imaginary number. true false i2 = sqrt-1 true false Complex numbers can be graphed on the xy coordinate plane. true false

OpenStudy (anonymous):

Complex numbers can be graphed on the xy coordinate plane. true

OpenStudy (jiteshmeghwal9):

i2=sqrt-1 is false

OpenStudy (jiteshmeghwal9):

isn't it @Yahoo! ?

OpenStudy (anonymous):

Any number in the form of a ± bi , where a and b are real numbers and b 0 is considered a pure imaginary number. false

OpenStudy (anonymous):

Alright! thanks!!

OpenStudy (anonymous):

@sue101 u shuld have written i^2

OpenStudy (jiteshmeghwal9):

ohhh ! i thought it \(2i\)

OpenStudy (anonymous):

oh sorry

OpenStudy (jiteshmeghwal9):

whyy ??

OpenStudy (anonymous):

\[i^2 = -1\]

OpenStudy (anonymous):

\[i = \sqrt{-1}\]

OpenStudy (anonymous):

Any number in the form of a ± bi , where a and b are real numbers and b 0 is considered a pure imaginary number. true (0 pts) false (1 pt) 1/1 point i2 = sqrt-1 true (0 pts) false (1 pt) 1/1 point Complex numbers can be graphed on the xy coordinate plane. true (0 pts) false (1 pt) 1/1 point